2. You can afford a $1150 per month mortgage payment. You've found a 30 year loan at 8% interest.

a) How big of a loan can you afford?

$

b) How much total money will you pay the loan company?

$

c) How much of that money is interest?

$
Answer #1

А B 1 a) $156,726.02 2 b) $414,000.00 3 c) $257,273.98

Calc:

A 1 a) 2 b) 3 C) B I=PV(8%/12,30*12,-1150) =1150*30*12 =B2-B1